# Blockchain Amplification Attack

### v2

@@ 3.2 Private monitoring node data
4 Modeling amplification
@@
4.1 Threat model
We first formalize an amplification attack on the P2P blockchain network (Figure 1). The attack leads modified nodes (i.e., victims) to forward invalid transactions, and increases their outgoing traffic usage and cost. The attacker attempts 1) to degrade the quality of the service provided by modified nodes, 2) to cause economic damage to modified nodes, and 3) to disrupt all nodes and users (especially high-frequency MEV bots or block builders that listen to the services provided by modified nodes). Beyond blockchain ecosystem participants, cloud service providers also potentially have an incentive to support attacks on modified nodes to increase traffic usage and, consequently, revenue.
@@ 4.1 Threat model
Figure 3: Invalid transaction creation
  1. 1.
    If the target modified does not check the account balance, the attacker can generate different addresses and send an infinite number of invalid transactions. The transaction cost is zero since there is no chance of inclusion.
  2. 2.
@@
In the first two cases, the attacker can specify an unreasonably high gas price to prioritize their transactions, without additional costs. The attacker can also combine multiple strategies (e.g., duplicating many invalid transactions using a past nonce).
Next, the attacker selects its target modified node. If the attacker receives any invalid transactions from a given peer, that peer is most likely the modified node. Alternatively, the attacker can monitor various centralized services’ txpools, identify the service(s) that accept(s) invalid transactions, and send $tx_{invalid}$ through RPC endpoints. The attacker finally sends $tx_{invalid}$ to one modified node, which accepts/inserts $tx_{invalid}$ to its txpool and forwards it with the rest of the network, including other modified nodes and regular nodes.
The modified nodes are tyipcally run by the same entity (as empirically observed in §6.2), they are most likely connected to minimize latencies. $tx_{invalid}$ would reach all the modified nodes. Even if some modified nodes belong to multiple entities, $tx_{invalid}$ would reach out to the rest as long as one of the nodes from each entity connects.
@@
4.2 Amplification factors
We want to precisely estimate the amplification factor to show how the attack scales up given the attacker’s input. The DoS/EDoS literature uses two metrics to quantify attack effectiveness: Traffic Amplification Factor (TAF) [33] and Economic Amplification Factor (EAF) [53], defined as
@@ 4.2 Amplification factors
$EAF=\frac{B_{victim}\cdot p_{victim}}{B_{attacker}\cdot p_{attacker}}=TAF\cdot\frac{p_{victim}}{p_{attacker}}\ ,$ (1)
@@
4.3 Network waste
To estimate how one invalid transaction amplifies in the network, we first derive 1) the number of bytes a regular node $i$ receives based on the number of peer connections (see blue arrows directed toward each regular node in Figure 1), 2) the number of bytes the entire network wastes based on the distributions of active connections on each peer (all the blue arrows). This calculation only considers the public nodes in a P2P network, but not the users (MEV bots and block miners/builders) who 1) privately connect to txpool providers or 2) listen to transactions through RPC endpoints outside the P2P network. The actual amplification factor would be the combination of 1) the amount of traffic that goes to users who listen to modified nodes and 2) TAF. We indeed notice that some modified nodes shut down the outbound connections to reduce outbound costs. Yet, they still incur the cost of propagating invalid transactions to users.
A regular node only hears $tx_{invalid}$ from modified nodes (not from other regular nodes). This allows us to focus solely on the connections from modified nodes to each regular node. Let the number of active connections for each node $i$ be $x_{i}$. The number of the modified node connection is $x_{i}\gamma$ where $\gamma$ is the ratio of modified nodes. This is based on the assumption that the node discovery process is random, thus the number of connected modified nodes increases linearly along with the number of connections by expectation. As described in §2, Ethereum nodes broadcast the transactions in two ways: 1) 0x02 message called “broadcast” ($a$ bytes), and 2) 0x08 message called “announcement” ($32$ bytes). Although Ethereum Wire protocol [15] suggests propagating 0x02 to a small number of peers for network efficiency, there is no clear consensus on the implementation, leading to variations across different software clients.
@@ 4.3 Network waste
Figure 4: Number of bytes a regular node $i$ receives based on the number of active connections.
As we already determine the waste for each regular node with a varying number of active connections, we proceed to calculate the total waste in the network by considering the distribution of peer connections, $g(x)$, which we empirically estimate in §6.1. There are $(1-\gamma)\cdot N=N_{\mathcal{R}}$ regular nodes where $N$ is the total number of nodes in the network. We multiply the amount of waste for each node ($f(x)$) and the distributions of peer connections $g(x)$ to get the overall network waste:
Network $\text{Network waste}=\sum_{i=1}^{N_{\mathcal{R}}}f(x_{i})g(x_{i})\ ,$ (2)
In our measurement, we estimate $g(x)$ for a part of the network (i.e., our connected peers), and smooth $g(x)$ into a continous form. In that case, the network waste is $N_{\mathcal{R}}\cdot\int_{0}^{1000}f(x)g(x)$ where we assume a maximum connections of 1,000 to ensure convergence and exclude an impractically large connections. We finally divide this by the transaction size ($a$) to get TAF. An adversary could set up modified nodes on their own to increase $\gamma$, but these nodes would also be attacked, making this strategy unproductive.
5 Empirical examples of attack
This section identifies the similar attack instances seen in the wild, and characterizes the size, intensity, cost, and strategies, and actors (the attacker – sender address, the victim – centralized services) involved in those attacks. While we cannot definitively pinpoint the motivations of those attackers, those attacks are executed in a manner consistent with our proposed attack (see the discussion in Appendix C).
5.1 Victims
We first attempt to find the instances of the attack in the txpool of centralized services as they are often the target of the attack. We find that some services deliver transactions much faster than others, but also propagate a significant number of dropped transactions. By comparing their txpools (from Flashbots’ data), we can practically figure out how fast, and how accurate their services are. For the latency comparison, we use the timestamp recorded in Flashbots’ data. For each on-chain transaction, we check how late each source propagates from the first moment Flashbots sees the transactions from any source. For accuracy, we calculate the ratio of the number of dropped transactions to the number of total transactions received. These numbers convey the quality of their propagation flow. Figure 5 (left) illustrates the bi-modal representation of latency vs. accuracy except for Flashbots’ “local” node. Each data point represents one week of data for each service. The $x$-axis is the median latency from all the transactions, whereas the $y$-axis is the ratio of dropped transactions. Services such as bloXroute, Chainbound, and Eden invest in infrastructure to optimize their propagation flow, resulting in potential latency reduction. Moreover, Flashbots’ database connects to bloXroute and Eden through gRPC, which is faster than the web socket commonly used by others (“delivery latency”). Additionally, a geographical advantage exists for services located close to the database—e.g., Flashbots’ local node likely operates a physical proxy to its database. The figure also suggests that those “fast” services fail to filter out transactions that are not supposed to persist in the network.
@@ 5.1 Victims
Figure 5: Comparing the centralized services’ txpools (left: accuracy-latency plot, right: dropped transactions between sources.
5.2 Identifying attack instances
@@
We find that some Ethereum accounts send an unexpectedly large number of transactions in a short range of time. We set two (conservative) thresholds to extrapolate automated non-human spamming behavior: if 1) we detect over 100 transactions from a single sender in any of the transaction pools within a single block interval (12 seconds) and 2) more than 95% of those transactions fail to be included in the chain. The second check is necessary to filter out a large entity (e.g., a popular gambling site) that submits many transactions at once without spamming intent.
Based on this heuristic, we find 2,591 instances from 345 Ethereum addresses from Flashbots data from September 1st, 2023 to January 11th, 2024. Figure 6 summarizes the number of attacks over time (on an hourly basis). There are more than 60 instances of attacks on November 26th between 7–8 AM (UTC). Many of them appear to send transactions from accounts with zero balances. We find similar cases on September 12th, 5 PM, October 7th, 5 PM, and November 20th, 9 PM. The number of attacks does not appear to be correlated with the Ether price fluctuation.
Figure 6: Number of attacks observed per hour, over time (September 1st, 2023–January 11th, 2024).
5.3 Characterizing attack behavior
@@
We first check the three types of invalid transactions discussed in §4.1 and Figure 3: 1) insufficient balance, 2) past nonce, and 3) duplicate transaction. To check the balance and nonce of the relevant accounts at the time of transaction submission, we use our archive node to fetch the state of the blockchain at the time of the timestamp. Given that there is a network delay between when the transaction enters a P2P network and when Flashbots records the timestamp, we look at the state of the account two blocks before the observation timestamp (i.e., at least 12 seconds in the past). Of 2,591 attack cases, we confirm that 536 instances (from 224 unique addresses) come from accounts that do not have a sufficient balance. Two services—bloXroute and Eden—have included those transactions during our observation window. In particular, some addresses do not have any transactions (send or receive) in their history whatsoever (i.e., zero Ether balance). In addition, three services (Chainbound, Eden, bloXroute) have accepted transactions that were using a past nonce (and thus, are invalid) during our measurement interval. Just like we identify invalid transactions from accounts with insufficient balance, we retrieve the last nonce of the corresponding address and identify 62 attacks from 27 unique addresses that use previously used nonce.
@@
Attackers also duplicate transactions in order to generate many invalid transactions. We categorize our identified attack instances based on the transaction parameters the attackers manipulate the most. There are four main parameters the attacker calibrates; 1) the gas limit, 2) the nonce, 3) the data field, and 4) the value of the transfer. This does not necessarily mean that the attackers only change those specific parameters, but sometimes they simultaneously permute several sets of parameters to increase the invalid transaction set, including other variables such as the recipient’s addresses, chain ID, and access list (introduced in EIP-2930). Table I summarizes the key statistics such as the transaction costs, intensity, and the size of the attack for each category. We first define key metrics: the size, cost, and intensity of the attack.
@@
About 63% of attack instances calibrate the gas limit parameter, but only two addresses are involved. Those two addresses manage to send 1,581 transactions with only 0.26 milliseconds between transactions on average. The largest attack carries 23MB of data between blocks (12 seconds). Since, on average, only one transaction lands on the chain, the transaction cost to the attacker is around 3 US dollars. Chainbound seems to be the main victim of this attack.
@@
To sum up, Table I illustrates each strategy’s pros and cons; the attacker can choose to manipulate different parameters depending on what the attacker wants to maximize/minimize: costs, size, intensity, and the number of victims (i.e., the centralized services in §5.1 that accept those invalid transactions).
TABLE I: Summary statistics for each attack type.
@@
6 Estimating model parameters
We next derive different sets of model parameters, which we then use to calculate amplification factors.
6.1 Peer connection distribution: $g(x)$
We formalize our method for estimating the number of active connections for each peer, $x_{i}$, and empirically computing its distribution, $g(x)$. Although studies on blockchain P2P network topology exist, the key novelty in our approach is to estimate active peer connections on the Ethereum mainnet without submitting test transactions (i.e., mediating ethical concerns and research costs). Instead, we reverse-engineer the number of peers based on propagation message types. We have 2.5 billion messages in five months (as explained in §3.2).
@@ 6.1 Peer connection distribution: $g(x)$
$l(\theta_{i},m_{2},m_{2}+m_{8})={m_{2}+m_{8}\choose m_{2}}\theta_{i}^{m_{2}}(1-\theta_{i})^{m_{8}}\ .$ (3)
@@
$\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}\Leftrightarrow\hat{x}_{i}=\left(\frac{1}{\hat{\theta}_{i}}\right)^{2}=\left(\frac{m_{2}+m_{8}}{m_{2}}\right)^{2}\ .$ (4)
@@
We apply Eqn. (4) for peers (of our nodes) that 1) use Geth or Erigon (before v2.49.0), and 2) we exclude nodes with a non-public git commit string, which is a sign of software modification as we discuss in the next section. If both of our monitoring nodes observe the same peer, we take the average between two. In total, we manage to estimate the number of peer connections for 6,005 nodes: mean and median of 41 and 31, respectively. For those, each peer has sent an average (median) of 293,121 (81,327) transaction messages. To make the distribution generalized/smooth and get $g(x)$, we apply the kernel density function (KDE) based on SciPy’s implementation & Scotts’ method for the bandwidth selection. Figure 7 shows the resulting PDF (Probability Density Function) and CDF (Cumulative Density Function) of active peer connections (green: estimate from Node A, orange: from Node B). The strong alignment between the estimations between two nodes indicates our method is robust. As an additional check, we ask one Ethereum node operator to provide us with the node name, node ID, and the maximum number of peers. The node uses an unmodified Geth node and sets the maximum number of peers to 40. Our estimator tells us that this node is expected to have $33.68$ peers. Given that not all the nodes are stable over time, our estimate appears to be consistent with this piece of ground truth.
Figure 7: Number of active peer connections for each regular node: left (PDF with a KDE smoothing), right (CDF).
6.2 Ratio of modified nodes: $\gamma$
We estimate the ratio of the modified nodes ($\gamma$) from the set of peers we connect. As described in §3, we capture transactions before entering the txpool, so we can empirically label the modified nodes that forward invalid transactions to us. First, we prepare three sets of invalid transactions as introduced in §4.1 and §5 to examine peers’ validation process: 1) insufficient balance, 2) past nonce, and 3) duplication (gas bump rule). We first label peers that have ever sent invalid transactions of the first two kinds: insufficient balance and past nonce. Our approach may mislabel modified nodes if their blockchain status is unsynced. However, the fact that the node forwards invalid transactions still allows attackers to conduct the same attack on unsynchronized nodes.

### v3

@@ 3.2. Private monitoring node data
Figure 2. An overview of our monitoring node.
4. Modeling amplification
@@
4.1. Threat model
We first formalize an amplification attack on the blockchain P2P network (Figure 1). The attack leads modified nodes (i.e., victims) to forward invalid transactions, and increases their outgoing traffic usage and cost. The attacker attempts 1) to degrade the quality of the service provided by modified nodes, 2) to cause economic damage to modified nodes, and 3) to disrupt all nodes and users (especially MEV bots or block builders that listen to the services provided by modified nodes). Beyond blockchain ecosystem participants, cloud service providers also potentially have an incentive to support attacks on modified nodes to increase traffic usage and, consequently, revenue.
@@ 4.1. Threat model
Figure 3. Invalid transaction creation
  1. (1)
    If the target modified node does not check the account balance, the attacker can generate new addresses and send an infinite number of invalid transactions. The transaction cost is zero since there is no chance of inclusion.
  2. (2)
@@
In the first two cases, the attacker can specify an unreasonably high gas price to prioritize their transactions in the txpool, without additional costs (because those transactions will be invalidated later). The attacker can also combine multiple strategies (e.g., duplicating many invalid transactions using a past nonce).
Next, the attacker selects its target modified node. If the attacker receives any invalid transactions from a given peer, that peer is most likely the modified node. Alternatively, the attacker can monitor various centralized entities’ txpools, identify the service(s) that accept(s) invalid transactions, and send $tx_{invalid}$ through RPC endpoints. The attacker finally sends $tx_{invalid}$ to one modified node, which accepts/inserts $tx_{invalid}$ to its txpool and forwards it with the rest of the network, including other modified nodes and regular nodes.
The modified nodes are tyipcally run by the same entity (as empirically observed in §6.2), they are most likely connected to minimize latencies. Hence, $tx_{invalid}$ would reach all the modified nodes. Even if some modified nodes belong to multiple entities, $tx_{invalid}$ would reach out to the rest as long as one of the nodes from each entity connects.
@@
4.2. Amplification factors
We want to precisely estimate the amplification factor to show how the attack scales up given the attacker’s input. The DoS/EDoS literature uses two metrics to quantify attack effectiveness: Traffic Amplification Factor (TAF) (Kumar, 2007) and Economic Amplification Factor (EAF) (Wang et al., 2016), defined as
@@ 4.2. Amplification factors
(1) $EAF=\frac{B_{victim}\cdot p_{victim}}{B_{attacker}\cdot p_{attacker}}=TAF\cdot\frac{p_{victim}}{p_{attacker}}\ ,$
@@
4.3. Network waste
To estimate how the original attack traffic amplifies in the network, we first derive 1) the number of bytes a regular node $i$ receives based on the number of peer connections (see blue arrows directed toward each regular node in Figure 1), 2) the number of bytes the entire network wastes based on the distributions of active connections on each peer (all the blue arrows). This calculation only considers the public nodes in a P2P network, but not the users (MEV bots and block miners/builders) who 1) privately connect to txpool providers or 2) listen to transactions through RPC endpoints outside the P2P network. The actual amplification factor would be the combination of 1) the amount of traffic that goes to users who listen to modified nodes and 2) TAF. We indeed notice that some modified nodes shut down the outbound connections to reduce outbound costs. Yet, they still incur the cost of propagating invalid transactions to users.
A regular node only hears $tx_{invalid}$ from modified nodes (not from other regular nodes). This allows us to focus solely on the connections from modified nodes to each regular node. Let the number of active connections for each regular node $i$ be $x_{i}$. The number of the modified node connection is $x_{i}\gamma$ where $\gamma$ is the ratio of modified nodes. This is based on the assumption that the node discovery process is random, thus the number of connected modified nodes increases linearly along with the number of connections by expectation. As described in §2, Ethereum nodes broadcast the transactions in two ways: 1) 0x02 message called “broadcast” ($a$ bytes), and 2) 0x08 message called “announcement” ($32$ bytes). Although Ethereum Wire protocol (Ethereum, 2023a) suggests propagating 0x02 to a small number of peers for network efficiency, there is no clear consensus on the implementation, leading to variations across different software clients.
@@ 4.3. Network waste
Figure 4. Number of bytes a regular node $i$ receives based on the number of active connections.
[…]
5. Empirical examples of attack
This section identifies the similar attack instances seen in the wild, and characterizes the size, intensity, cost, and strategies, and actors (the attacker – sender address, the victim – centralized entities) involved in those attacks. While we cannot definitively pinpoint the motivations of those attackers, those attacks are executed in a manner consistent with our proposed attack (see the discussion in Appendix C).
5.1. Victims
We first attempt to find the instances of the attack in the txpool of centralized entities as they are often the target of the attack. We find that some services deliver transactions much faster than others, but also propagate a significant number of dropped transactions. By comparing their txpools (from Flashbots’ data), we can practically figure out how fast, and how accurate their services are. For the latency comparison, we use the timestamp recorded in Flashbots’ data. For each on-chain (finalized) transaction, we check how late each source propagates from the first moment Flashbots sees the transactions from any source. For accuracy, we calculate the ratio of the number of dropped transactions to the number of total transactions received. These numbers convey the quality of their propagation flow. Figure 5 (left) illustrates the bi-modal representation of latency vs. accuracy except for Flashbots’ “local” node. Each data point represents one week of data for each service. The $x$-axis is the median latency from all the transactions, whereas the $y$-axis is the ratio of dropped transactions. Services such as bloXroute, Chainbound, and Eden invest in infrastructure to optimize their propagation flow, resulting in potential latency reduction. Moreover, Flashbots’ database connects to bloXroute and Eden through gRPC, which is faster than the web socket commonly used by others (“delivery latency”). Additionally, a geographical advantage exists for services located close to the database—e.g., Flashbots’ local node likely operates a physical proxy to its database. The figure also suggests that those “fast” services fail to filter out transactions that are not supposed to persist in the network.
@@ 5.1. Victims
Figure 5. Comparing the centralized entities’ txpools (left: accuracy-latency plot, right: dropped transactions between sources.
5.2. Identifying and characterizing attack behavior
@@
We find that some Ethereum accounts send an unexpectedly large number of transactions in a short range of time. We set two (conservative) thresholds to extrapolate automated non-human spamming behavior (as a lower bound): if 1) we detect over 100 transactions from a single sender within a single block interval (12 seconds) and 2) more than 95% of those transactions fail to be included in the chain. The second check is necessary to filter out a large entity (e.g., a popular gambling site) that submits many transactions at once without spamming intent. Based on this heuristic, we find 2,591 instances from 345 Ethereum addresses from Flashbots data from September 1st, 2023 to January 11th, 2024.
@@
We next characterize attack like behavior, in particular, three types of invalid (attack) transactions discussed in §4.1 and Figure 3: 1) insufficient balance, 2) past nonce, and 3) duplicate transaction. To check the account balance and nonce at the time of transaction submission, we use our archive node to fetch the state of the blockchain at the time of the timestamp. Given a network delay between when the transaction enters a P2P network and when Flashbots records the timestamp, we look at the state of the account two blocks before the observation timestamp (i.e., at least 12 seconds in the past). Of 2,591 attack cases, we confirm that 536 instances (from 224 unique addresses) come from accounts that do not have a sufficient balance. Two services (bloXroute and Eden) have included those transactions during our observation window. In particular, some addresses do not have any transactions (send or receive) in their history whatsoever (i.e., zero Ether balance). In addition, three services (Chainbound, Eden, bloXroute) have accepted transactions that are using a past nonce (and thus, are invalid) during our measurement interval. Just like we identify invalid transactions from accounts with insufficient balance, we retrieve the last nonce of the corresponding address and identify 62 attacks from 27 unique addresses that use previously used nonce.
@@
Attackers also duplicate transactions in order to generate many invalid transactions. We categorize our identified attack instances based on the transaction parameters the attackers manipulate the most. There are four main parameters the attacker calibrates; 1) the gas limit, 2) the nonce, 3) the data field, and 4) the transfer value. This does not necessarily mean that the attackers only change one specific parameter, but sometimes they simultaneously permute several sets of parameters to increase the invalid transaction set, including other variables such as the recipient’s addresses, chain ID, and access list (introduced in EIP-2930). Table 1 summarizes the key statistics such as the transaction costs, intensity, and the size of the attack for each category. We first define key metrics: the size, cost, and intensity of the attack.
@@
About 63% of attack instances calibrate the gas limit parameter, but only two addresses are involved. Those two addresses manage to send 1,581 transactions with only 0.26 milliseconds between transactions on average. The largest attack carries 23MB of unique data between blocks (12 seconds). Since, on average, only one transaction lands on the chain, the transaction cost to the attacker is around 3 US dollars. Chainbound seems to be the main victim of this attack.
@@
To sum up, Table 1 illustrates each strategy’s pros and cons; the attacker can manipulate different parameters to prepare many invalid transactions, depending on what the attacker wants to maximize/minimize: costs, size, intensity, and the number of victims (i.e., the centralized entities).
@@
Table 1. Summary statistics for each attack type.
6. Estimating model parameters
We next derive our model parameters, which we then use to calculate amplification factors.
6.1. Peer connection distribution: $g(x)$
We formalize our method for estimating the number of active connections for each peer, $x_{i}$, and empirically computing its distribution, $g(x)$. Although studies on blockchain P2P network topology exist, the key novelty in our approach is to estimate active peer connections on the Ethereum mainnet without submitting test transactions (i.e., mediating ethical concerns and research costs). Instead, we reverse-engineer the number of peers based on propagation message types. We collect 2.5 billion messages in five months (as explained in §3.2) to perform this method.
@@ 6.1. Peer connection distribution: $g(x)$
(2) $l(\theta_{i},m_{2},m_{2}+m_{8})={m_{2}+m_{8}\choose m_{2}}\theta_{i}^{m_{2}}(1-\theta_{i})^{m_{8}}\ .$
@@
(3) $\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}\Leftrightarrow\hat{x}_{i}=\left(\frac{1}{\hat{\theta}_{i}}\right)^{2}=\left(\frac{m_{2}+m_{8}}{m_{2}}\right)^{2}\ .$
@@
We apply Eqn. (3) for peers (of our nodes) that 1) use Geth or Erigon (before v2.49.0), and 2) we exclude nodes with a non-public git commit string, which is a sign of software as we discuss in §6.2. If our two monitoring nodes observe the same peer, we take the average between two. In total, we manage to estimate the number of peer connections for 6,005 nodes: mean and median of 41 and 31, respectively. For those, each peer has sent an average (median) of 293,121 (81,327) transaction messages. To make the distribution generalized/smooth and get $g(x)$, we apply the kernel density function (KDE) on SciPy’s implementation & Scotts’ method for the bandwidth selection. Figure 6 shows the resulting PDF (Probability Density Function) and CDF (Cumulative Density Function) of active peer connections (green: estimate from Node A, orange: from Node B). The strong alignment between the two estimates indicates our method is robust. As an additional check, we ask one Ethereum node operator to provide us with the node name, node ID, and the maximum number of peers. The node uses an unmodified Geth node and sets the maximum number of peers to 40. Our estimator tells us that this node is expected to have $33.68$ peers. Given that not all the nodes are active/stable over time, our estimate appears to be consistent with this piece of ground truth.
Figure 6. Number of active peer connections for each regular node: left (PDF with a KDE smoothing), right (CDF).
6.2. Ratio of modified nodes: $\gamma$
We estimate the ratio of the modified nodes ($\gamma$) from the set of peers we connect. As described in §3.2, we capture transactions before entering the txpool, so we can empirically label the modified nodes that forward invalid transactions to us. First, we prepare three sets of invalid transactions as introduced in §4.1 and §5 to examine peers’ validation process: 1) insufficient balance, 2) past nonce, and 3) duplication (gas bump rule). We first label peers that have ever sent invalid transactions of the first two kinds: insufficient balance and past nonce. Our approach may mislabel modified nodes if their blockchain status is unsynced. However, the fact that the node forwards invalid transactions still allows attackers to conduct the same attack on unsynchronized nodes.